A work of 657 J is done on a system that releases 434 J of heat. What is the energy change in the system?


1
Expert's answer
2022-04-20T11:20:03-0400

U = Q + W

When work is done on the system it is positive. When heat is released it's negative.


U = 657 - 434 = 223 J